Allison:

If MK= 10 m, Find the length of arc MKL

Allison:

1 attachment
dude:

This is spicier We want to find the red part first but in order to find the red we need to find the rest of the blue There is a right angle (I marked blue) They gave us 52 degrees 90- 52 = 38 Now from there do 180 - 38 to find greek letter Your r is MK divided by 2

1 attachment
Allison:

But wait, the next one gets hotter hahaah

dude:

D:

Allison:

What's MK? 52?

dude:

`If MK= 10 m, Find the length of arc MKL`

Allison:

r = 5

dude:

Yes

Allison:

Greek letter is 38

dude:

No 180 - 38

Allison:

142

dude:

Yes

Allison:

2 times pi times 5 times 142/360

dude:

Yes

Allison:

12.39

dude:

Yes

dude:

Fixing for future reference Asks for MKL, not ML Do 360 - 142 instead \(s= 2\pi (5)(\dfrac{218}{360})\) This should give you 19.02